OpenStudy (anonymous):

Given the integral of (1/(1+9x^2)dx taking u =3x. Use integration by substitution Question is can someone explain to me how they got u as 3x.As for me would suggest u=1+9x^2

OpenStudy (bahrom7893):

let u = 3x du = 3dx The integral can be rewritten as: (1/3)*(3dx/(1+9x^2))

OpenStudy (anonymous):

but hwy 3x

OpenStudy (bahrom7893):

Now if we substitute u = 3x, du = 3dx, we get: (1/3)*(du/(1+u^2) = (1/3) (1/(1+u^2)*du

OpenStudy (bahrom7893):

Because the bottom seems to be in the form of 1+a^2, where a^2 = 9x^2 = (3x)^2

OpenStudy (bahrom7893):

So after you substitute u: (1/3) (1/(1+u^2)*du, notice that the integral is just inverse tangent

OpenStudy (bahrom7893):

So you can leave the constant alone, and if you integrate (1/(1+u^2))*du, you'll get ArcTan(u)

OpenStudy (bahrom7893):

Now just multiply by that constant outside to get (1/3)*ArcTan(u), but u're not done yet: u = 3x, so now we have: (1/3)*Arctan(3x) + C (don't forget the arbitrary constant, this is an indefinite integral).